OpenStudy (anonymous):

the product of 7/10 with the quotient of 3/2 and 7/11?

OpenStudy (luffingsails):

Sorry, missed this one So from verbal to formula, they are asking for \[\frac{7}{10} * \frac{\frac{3}{2}}{\frac{7}{11}}\] After simplification on the right \[\frac{7}{10} * \frac{3}{2} * \frac{11}{7}\] Just multiply out and simplify afterwards.

OpenStudy (anonymous):

\[\frac{7}{10} *\left( \left(\frac{3}{2}\right) /\left( \frac{7}{11}\right)\right)=\frac{33}{20} \]